# CROSS-GAiT: Cross-Attention-Based Multimodal Representation Fusion for Parametric Gait Adaptation in Complex Terrains

### v2

@@ Abstract
We present CROSS-GAiT, a novel algorithm for quadruped robots that uses Cross Attention to fuse terrain representations derived from visual and time-series inputs, including linear accelerations, angular velocities, and joint efforts. These fused representations are used to adjust the robot’s step height and hip splay, enabling adaptive gaits that respond dynamically to varying terrain conditions. We generate these terrain representations by processing visual inputs through a masked Vision Transformer (ViT) encoder and time-series data through a dilated causal convolutional encoder. The cross-attention mechanism then selects and integrates the most relevant features from each modality, combining terrain characteristics with robot dynamics for better informed gait adjustments. CROSS-GAiT uses the combined representation to dynamically adjust gait parameters in response to varying and unpredictable terrains. We train CROSS-GAiT on data from diverse terrains, including asphalt, concrete, brick pavements, grass, dense vegetation, pebbles, gravel, and sand. Our algorithm generalizes well and adapts to unseen environmental conditions, enhancing real-time navigation performance. CROSS-GAiT was implemented on a Ghost Robotics Vision 60 robot and extensively tested in complex terrains with high vegetation density, uneven/unstable surfaces, sand banks, deformable substrates, etc. We observe at least a 7.04% reduction in IMU energy density and a 27.3% reduction in total joint effort, which directly correlates with increased stability and reduced energy usage when compared to state-of-the-art methods. Furthermore, CROSS-GAiT demonstrates at least a 64.5% increase in success rate and a 4.91% reduction in time to reach the goal in four complex scenarios. Additionally, the learned representations perform 4.48% better than the state-of-the-art on a terrain classification task.
@@ I Introduction
Recent advancements in quadruped robot platforms have been used for autonomous navigation in challenging outdoor environments, such as rocky surfaces, deformable terrains, and densely vegetated areas where wheeled or tracked robots do not work well [1, 2, 3, 4, 5, 6]. To achieve consistent performance, quadruped robots must rely on accurate terrain understanding to navigate efficiently, which is essential for maintaining stability, reducing energy consumption, and supporting long missions in complex environments [7, 8, 9].
Figure 1: Comparison of CROSS-GAiT with other state-of-the-art methods while navigating a Ghost Robotics Vision 60 robot through a complex scenario that contains a paved concrete path, thick bushes, short grass, and pebbles: CROSS-GAiT adapts gaits parametrically. CROSS-GAiT outperforms all other compared methods [10, 11, 12] based on the navigation metrics defined in Section V-C. We observe significant improvement in the success rate. The two graphs represent IMU joint effort and IMU energy density (both median-filtered to reduce noise artifacts) over time, illustrating the reduced energy consumption and improved stability achieved by CROSS-GAiT.
Exteroceptive sensors, such as cameras and LiDARs, play a key role in providing the terrain understanding needed for navigating complex, unstructured environments. Numerous approaches leverage these sensors [13, 12, 14, 4], combined with learning-based methods and multi-modal fusion techniques, to perform terrain segmentation, classification, and generate occupancy maps for reliable traversability estimates [15, 12, 16, 17]. However, current methods mostly rely on manually annotated terrain datasets, which are labor-intensive, time-consuming, and prone to human bias.
@@
Effective gait adaptation is critical in complex scenarios where avoiding unstable regions is infeasible, such as when a quadruped robot encounters dense vegetation, sand banks, or rocky terrains. Using a fixed gait in such conditions may lead to increased energy consumption and instability [26, 27, 7, 8]. Several methods have been proposed to switch between pre-defined gaits based on estimated traversability calculated using simple rules [10, 11], but their application has been limited.
Main contributions: To address these limitations, we present CROSS-GAiT (Cross-Attention-Based Multi-modal Representation Fusion for Gait Adaptation in Complex Terrains), a novel approach that fuses terrain representations from visual, IMU, and proprioceptive inputs using a cross-attention transformer network. Our fused representation is leveraged to dynamically adjust gait parameters that adapt to the terrain, reduce energy consumption and increase perceived stability during navigation. The novel components of our work include:
  • A novel fusion algorithm: Fuses latent representations of visual data with time-series data (linear accelerations, angular velocities, and proprioception) using cross-attention, corresponding to a comprehensive latent representation of the environment. Trained with supervised contrastive loss, CROSS-GAiT achieves 98.45% terrain classification accuracy when evaluated on test data from the dataset described in Section V-B, outperforming state-of-the-art methods by 4.48% in terrain classification using the learned representations.
  • A parametric gait adaptation algorithm: Adjusts step height and hip splay to generate custom gaits beyond default options available on quadruped platforms. Our algorithm learns gait parameters that improve stability by 7.04% and reduce joint effort by 27.3%, compared to existing methods.
  • A real-time implementation: Our system runs at 60 Hz on an Intel NUC 11 with an Intel i7 CPU and NVIDIA RTX 2060 GPU. We demonstrate CROSS-GAiT on a Ghost Robotics Vision60, evaluating its performance in four complex scenarios on top of an exteroceptive navigation algorithm [4].
@@ II-A Perception Methods for Terrain Understanding
Proprioceptive sensory data, such as joint forces and inertial measurements, are widely used for instability detection during navigation [18, 11]. Dey et al. [18] predict slippage using proprioceptive time series, while Kumar et al. [19] propose a standalone proprioceptive reinforcement learning-based navigation algorithm for outdoor terrains. However, proprioceptive methods struggle to anticipate future terrains, limiting their utility [26]. To overcome this, studies have fused proprioceptive data with LiDAR or visual inputs [10].
Figure 2: Overall architecture of CROSS-GAiT: Our method utilizes a vision transformer-based masked autoencoder for image data and a dilated causal convolutional encoder for IMU and joint effort data to generate latent representations. These are combined using a cross-attention transformer network. The combined latent representation is then passed through a multilayer perceptron-based regressor to obtain the final output: a set of gait parameters, including step height and hip splay. A dynamic window mechanism is applied to ensure smooth gait transitions during terrain navigation. The overall system runs at 60 Hz on the implemented system.
@@ II-B Gait Change Mechanisms for Navigation in Complex Environments
Adaptive gait strategies help maintain stability and reduce energy consumption in complex environments [26, 27]. Several studies have examined switching between predefined gaits for stability and energy efficiency [10, 11].
@@ II-C Cross-Attention Multi-modal Fusion
Cross-attention mechanisms have demonstrated significant success in fusing information from multiple modalities across various domains for useful downstream tasks such as object detection [36], image and sentence matching [37], information fusion [38], emotion detection [39], recommendor systems [40] and soft robot manipulation [41]. For example, BronchoCopilot, a multi-modal reinforcement learning agent designed for autonomous robotic bronchoscopy, employs cross-attention mechanisms to fuse visual information from the bronchoscope camera with the robot’s estimated pose. This fusion enables the system to navigate complex airway environments more effectively. Unlike these methods, we fuse proprioceptive time-series data with exteroceptive visual data to generate a comprehensive terrain representation, enabling dynamic gait parameter adjustments for quadruped robots navigating diverse terrains.
@@ III-A Notations
We use the following notations: $N$ for architectural counts (e.g., layers or blocks), $n$ and $m$ for input and batch sizes, $J$ for joint effort data, $I$ for image inputs, $U$ for IMU data, and $Z$ for latent representations. $\mathcal{L}$ represents losses, $W$ denotes weight matrices, and $Q$, $K$, and $V$ are query, key, and value vectors in attention mechanisms. $d$ refers to dimensions, $t$ to time, $g$ to gait parameters, and $\theta$ to terrains.
@@ III-E Cross-Attention Transformer Mechanism
Cross-attention fuses information from multiple modalities by allowing one modality to attend to relevant features from another. In multimodal learning, it enables the model to focus on the most informative aspects of each modality, such as visual and time-series inputs [38, 47]. Queries (Q) from one modality attend to keys (K) and values (V) from another, integrating complementary information for richer representations. Other components of the transformer, like add & norm layers, residual connections, and feedforward networks, remain unchanged [48].
@@ IV Our Approach: CROSS-GAiT
We present CROSS-GAiT, a method for navigating complex outdoor environments inspired by how humans adapt their gait using visual and proprioceptive feedback. Similar to humans adjusting their gait for stability on diverse terrains, CROSS-GAiT promotes stable and energy-efficient gaits for surfaces like grass, sand, asphalt, and rocky terrains. Our approach focuses on four key aspects: generating latent representations from visual data, generating representations from time-series data, fusing these multi-modal inputs via cross-attention, and using the fused representation to generate adaptive gait parameters. The overall architecture is outlined in Figure 2.
@@ IV-A Latent Representations of Visual Data
We initialize the model with pre-trained weights from [43] and fine-tune it on our custom dataset (described in Section V-B), ensuring that the learned representations are well-suited to the terrains the robot will encounter.
@@ IV-C Combining Latent Representations of Visual Data and Time Series Data
We pass the result of the cross-attention operation (3), $\mathbf{Z}^{t}_{\text{imd}}\in\mathbb{R}^{1\times d_{v}}$, through the rest of the cross-attention block, including the add & norm layer, and a feedforward network, which introduces nonlinearity and refines the fused representation. This output is then fed into $N_{s}a$ self-attention transformer blocks to produce the final output $\mathbf{Z}^{t}_{\text{comb}}\in\mathbb{R}^{1\times d_{v}}$. The cross-attention fuses relevant features from both modalities, while the self-attention blocks capture deeper dependencies and interactions between the image and time-series data.
@@
Figure 3: Image reconstruction outputs from the masked autoencoder (MAE) trained for feature encoding. 75% of the input images are masked and fed into the MAE for reconstruction. The trained MAE captures fine-grained details such as dry leaves and the granularity of sand/pebbles. During inference, only the encoder is used to obtain latent representations that encode image features for complex terrains.
@@
We define a gait parameter cost function (5) to identify parameter combinations that minimize energy consumption and increase stability for the terrains in the training dataset (Section V-B). The parameter pair with the lowest cost for each terrain is selected to generate the gait parameter labels used for training. In this cost function, $\tilde{\omega}_{t}$, $\tilde{a}_{z}$, $\tilde{J}_{\text{effort}}$, $\alpha$, $\beta$, and $\gamma$ represent the normalized angular velocity, linear acceleration along the z axis, r.m.s joint effort, and their corresponding weights. Parameter selection for training is detailed in Section V-B.
@@ IV-D Model for Gait Parameter Generation
We pass the combined latent representation $\mathbf{Z}^{t}_{\text{comb}}$ through a multi-layer perceptron (MLP) regressor to predict suitable gait parameters ($\mathbf{g}$), specifically, the robot’s step height and hip splay. We train the regressor using the MSE loss, based on the training labels generated using the analysis of (5) on the dataset. During this phase, all layers preceding the regressor, including the Masked Autoencoder, Dilated Causal Convolutional Encoder, and cross-attention/sel-attention layers, are kept frozen.
The regressor contains two fully connected hidden layers with ReLU activation, while the output layer does not use any activation function. ReLU introduces non-linearity, enabling the network to model complex relationships between the combined representation and gait parameters.
The model learns continuous mappings from sensory inputs to gait parameters, allowing it to generalize to unseen terrains by interpolating or extrapolating from training patterns. This approach ensures dynamic gait adaptation based on real-time sensory data, providing a robust solution for navigating complex terrains. A dynamic windowed approach smooths consecutive gait parameter predictions, ensuring stable and controlled gait transitions.
@@ V Analysis and Results
Figure 4: [TOP]: Illustration of the cost values evaluated for different gait parameter pairs using the cost function defined in Equation 5. The circled parameter pairs represent the selected labels for training. [BOTTOM]: Sample terrain images corresponding to each surface type in the above plots. Paramater pairs that lead to the lowest cost (dark blue) in each terrain are highlighted in green circles.
@@ V-A Implementation
We implemented CROSS-GAiT using PyTorch and trained all models on an Nvidia A5000 GPU. For real-world experiments, we used the Ghost Vision 60 robot from Ghost Robotics, equipped with a front-facing wide-angle camera and an onboard Intel NUC 11 (Intel i7 CPU, NVIDIA RTX 2060 GPU). Our system operates in real-time at 60Hz on this hardware configuration. Path planning was implemented using [4]. We sample a 6-channel IMU and joint effort data, comprising 12 channels (one for each motor, measured in Nm), at 25 Hz. We combine these data streams to form the input time series data $\mathbf{X}^{t}_{\text{TS}}$, resulting in $n_{\text{ch}}=18$. $d_{\text{TS}}=100$, is chosen empirically to capture the frequency components of the robot’s cyclic gaits. $d_{m}=160$ based on [46].
@@
Figure 5: Robot trajectories in unstructured terrains using CROSS-GAiT and comparison methods. Joint Effort (left) and IMU Energy Density (right) plots (median-filtered to reduce noise) are shown above each scenario. CROSS-GAiT reaches the goal faster with lower joint effort, while maintaining comparable IMU energy density to other methods.

### v3

@@ Abstract
We present CROSS-GAiT, a novel algorithm for quadruped robots that uses Cross Attention to fuse terrain representations derived from visual and time-series inputs; including linear accelerations, angular velocities, and joint efforts. These fused representations are used to continuously adjust two critical gait parameters (step height and hip splay), enabling adaptive gaits that respond dynamically to varying terrain conditions. To generate terrain representations, we process visual inputs through a masked Vision Transformer (ViT) encoder and time-series data through a dilated causal convolutional encoder. The Cross Attention mechanism then selects and integrates the most relevant features from each modality, combining terrain characteristics with robot dynamics for informed gait adaptation. This fused representation allows CROSS-GAiT to continuously adjust gait parameters in response to unpredictable terrain conditions in real-time. We train CROSS-GAiT on a diverse set of terrains including asphalt, concrete, brick pavements, grass, dense vegetation, pebbles, gravel, and sand and validate its generalization ability on unseen environments. Our hardware implementation on the Ghost Robotics Vision 60 demonstrates superior performance in challenging terrains, such as high-density vegetation, unstable surfaces, sandbanks, and deformable substrates. We observe at least a 7.04% reduction in IMU energy density and a 27.3% reduction in total joint effort, which directly correlates with increased stability and reduced energy usage when compared to state-of-the-art methods. Furthermore, CROSS-GAiT demonstrates at least a 64.5% increase in success rate and a 4.91% reduction in time to reach the goal in four complex scenarios. Additionally, the learned representations perform 4.48% better than the state-of-the-art on a terrain classification task.
@@ I Introduction
Recent advancements in quadruped robot platforms have been used for autonomous navigation in challenging outdoor environments, such as rocky surfaces, deformable terrains, and densely vegetated areas where wheeled or tracked robots do not work well [1, 2, 3, 4, 5, 6]. To achieve consistent performance and to navigate efficiently, quadruped robots must rely on accurate terrain understanding, which is essential for maintaining stability, reducing energy consumption, and supporting long missions in complex environments [7, 8, 9].
Figure 1: Comparison of CROSS-GAiT with state-of-the-art methods while navigating a Ghost Robotics Vision 60 robot through a complex scenario that contains a paved concrete path, thick bushes, grass, and pebbles: CROSS-GAiT continuously adapts gait parameters (step height and hip splay) based on multi-sensory observations and outperforms all other comparison methods [10, 11, 12] in terms of success rate and time to reach the goal. [TOP] two graphs represent IMU joint effort and IMU energy density (both median-filtered to reduce noise artifacts) over time, illustrating the reduced energy consumption and improved stability achieved by CROSS-GAiT in comparison to the other methods. [BOTTOM] two graphs present the continuous variation of gait parameters according to different terrain conditions. For instance, CROSS-GAiT uses high steps in dense vegetation to reduce leg entrapment while increasing hip splay during instabilities or sinkage, even when the camera is occluded by tall grass.
Exteroceptive sensors, such as cameras and LiDARs, play a key role in providing the terrain understanding needed for navigating such unstructured environments. Numerous approaches leverage these sensors [13, 12, 14, 4], combined with learning-based methods and multi-modal fusion techniques, to perform terrain segmentation, classification, and generate occupancy maps for reliable traversability estimates [15, 12, 16, 17]. However, current methods mostly rely on manually annotated terrain datasets, which are labor-intensive, time-consuming, and prone to human bias.
@@
Effective gait adaptation is critical in complex scenarios where avoiding unstable regions is infeasible, such as when a quadruped robot encounters dense vegetation, sandy or rocky terrains. Using a fixed gait in such conditions may lead to increased energy consumption and instability [26, 27, 7, 8]. Several methods have been proposed to switch between pre-defined gaits based on estimated traversability calculated using simple rules [10, 11], but their application has been limited.
Main contributions: To address these limitations, we present CROSS-GAiT (Cross-Attention-Based Multi-modal Representation Fusion for parametric Gait Adaptation in Complex Terrains), a novel approach that fuses terrain representations from visual, IMU, and proprioceptive inputs using a cross-attention transformer network. Our fused representation is leveraged to continuously adjust gait parameters that adapt to the terrain, reduce energy consumption and increase perceived stability during navigation. The novel components of our work include:
  • A novel multi-modal fusion algorithm that fuses latent representations of visual data with time-series data (linear accelerations, angular velocities, and proprioception) using cross-attention, resulting in a comprehensive latent representation of the environment. Trained with supervised contrastive loss, CROSS-GAiT achieves 98.45% terrain classification accuracy when evaluated on test data from the dataset described in Section V-B, outperforming state-of-the-art and classical MLP-based fusion methods at least by 3.18% in terms of terrain classification accuracy using the learned representations.
  • A novel parametric gait adaptation algorithm that continuously adjusts two critical gait parameters (step height and hip splay) enabling smooth and dynamic transitions in the quadruped robot’s gaits. Unlike traditional methods that switch between a predefined set of discrete gaits, our approach allows for real-time and continuous gait modulation based on environmental conditions and proprioceptive feedback. This continuous adaptation improves stability by 7.04% and reduces joint effort by 27.3% compared to existing gait adaptation techniques.
  • A real-time implementation: Our model is designed to be lightweight for real-time execution on edge hardware. CROSS-GAiT achieves a processing rate of approximately 60 Hz on a mobile computing platform equipped with an Intel NUC 11 (Intel i7 CPU, NVIDIA RTX 2060 GPU). We validate CROSS-GAiT on the Ghost Robotics Vision60 quadruped, demonstrating its effectiveness in complex outdoor environments using an exteroceptive navigation algorithm [4], showcasing robust performance in real-world deployment scenarios.
@@ II-A Perception Methods for Terrain Understanding
Proprioceptive sensory data, such as joint forces and inertial measurements, are widely used for instability detection during navigation [18, 11]. Dey et al. [18] predict slippage using time-series proprioception data, while Kumar et al. [19] propose a standalone proprioceptive reinforcement learning-based navigation algorithm for outdoor terrains. However, proprioceptive methods struggle to anticipate future terrains, limiting their utility [26]. To overcome this, studies have fused proprioceptive data with LiDAR or visual inputs [10].
Figure 2: Overall architecture of CROSS-GAiT: Our method utilizes a vision transformer-based masked autoencoder for image data and a dilated causal convolutional encoder for IMU and joint effort data to generate latent representations. These are combined using a cross-attention transformer network. The combined latent representation is then passed through a multilayer perceptron-based regressor to obtain the final output: a set of gait parameters, including step height and hip splay. A dynamic window mechanism is applied to ensure smooth gait transitions during terrain navigation. The overall system runs around 60 Hz on the implemented system.
@@ II-B Gait Change Mechanisms for Navigation in Complex Environments
Gait adaptation helps maintain quadraped stability and reduce energy consumption in complex environments [26, 27]. Several studies have examined switching between predefined gaits for stability and energy efficiency [10, 11].
@@ II-C Cross-Attention Multi-modal Fusion
Cross-attention mechanisms have demonstrated significant success in fusing information from multiple modalities across various domains for useful downstream tasks such as object detection [36], image and sentence matching [37], information fusion [38], emotion detection [39], recommendor systems [40] and soft robot manipulation [41]. For example, BronchoCopilot, a multi-modal reinforcement learning agent designed for autonomous robotic bronchoscopy, employs cross-attention mechanisms to fuse visual information from the bronchoscope camera with the robot’s estimated pose. This fusion enables the system to navigate complex airway environments more effectively.
In contrast to these existing methods, we fuse proprioceptive time-series data with exteroceptive visual data to generate a comprehensive terrain representation, enabling dynamic gait parameter adjustments for quadruped robots navigating diverse terrains.
@@ III-A Notations
We use the following notations: $N$ for architectural counts (e.g., layers or blocks), $n$ and $m$ for input and batch sizes, respectively, $J$ for joint effort data, $I$ for image inputs, $U$ for IMU data, and $Z$ for latent representations. $\mathcal{L}$ represents losses, $W$ denotes weight matrices, and $Q$, $K$, $V$ are query, key, value vectors in attention mechanisms, respectively. $d$ refers to dimensions, $t$ to time, $g$ to gait parameters, and $\theta$ to terrains.
@@ III-E Cross-Attention Transformer Mechanism
Cross-attention fuses information from multiple modalities by allowing one modality to attend to relevant features from another. This allows a model to selectively emphasize relevant features from each modality, reducing redundancy and enhancing contextual understanding. In multimodal learning, it enables the model to focus on the most informative aspects of each modality, such as visual and time-series inputs [38, 47]. Queries, Q from one modality attend to keys K and values V from another, integrating complementary information for richer representations. Other components of the transformer, like add & norm layers, residual connections, and feedforward networks, remain unchanged [48]. Overall, unlike concatenation followed by MLPs [6], which apply static transformations, cross-attention enables adaptive feature alignment, leading to more robust and informative representations.
@@ IV Our Approach: CROSS-GAiT
We present CROSS-GAiT, a method for navigating complex outdoor environments inspired by how humans adapt their gait using visual and proprioceptive feedback. Similar to humans adjusting their gait for stability on diverse terrains, CROSS-GAiT ensures stable and energy-efficient gaits for surfaces like grass, sand, asphalt, and rocky terrains. Our approach focuses on four key aspects: generating latent representations from visual data, generating representations from time-series data, fusing these multi-modal inputs via cross-attention, and using the fused representation to generate adaptive gait parameters. The overall architecture is outlined in Figure 2.
@@ IV-A Latent Representations of Visual Data
We initialize the model with pre-trained weights from [43] and fine-tune it on our custom dataset (described in Section V-B), ensuring that the learned representations are well-suited to the terrains the robot would encounter.
@@ IV-B Latent Representation of Time Series Data
Figure 3: Image reconstruction outputs from the masked autoencoder (MAE) trained for feature encoding. 75% of the input images are masked and fed into the MAE for reconstruction. The fine-tuned MAE captures fine-grained details such as dry leaves and the granularity of sand/pebbles. During inference, only the encoder is used to obtain latent representations that encode image features for complex terrains.
@@ IV-C Combining Latent Representations of Visual Data and Time Series Data
We pass the result of the cross-attention operation (3), $\mathbf{Z}^{t}_{\text{imd}}\in\mathbb{R}^{1\times d_{v}}$, through the rest of the cross-attention block, including the add & norm layer, and a feedforward network, which introduces nonlinearity and refines the fused representation. This output is fed into $N_{sa}$ self-attention transformer blocks to produce the final output $\mathbf{Z}^{t}_{\text{comb}}\in\mathbb{R}^{1\times d_{v}}$. The cross-attention fuses relevant features from both modalities, while the self-attention blocks capture deeper dependencies and interactions between the image and time-series data.
@@
We define a gait parameter cost function (5) to identify parameter combinations that minimize energy consumption and increase stability for the terrains in the training dataset (Section V-B). The parameter pair with the lowest cost for each terrain is selected to generate the gait parameter labels used for training. In this cost function, $\tilde{\omega}_{t}$, $\tilde{a}_{z}$, $\tilde{J}_{\text{effort}}$, $\alpha$, $\beta$, and $\gamma$ represent the normalized angular velocity, linear acceleration along the z axis, RMS joint effort, and their corresponding weights. Parameter selection for training is detailed in Section V-B.
Figure 4: [TOP]: Illustration of the cost values evaluated for different gait parameter pairs using the cost function defined in Equation 5. The x axis denotes step height and the y axis represents hip splay. The circled parameter pairs represent the selected labels for training the regression. [BOTTOM]: Sample terrain images corresponding to each surface type in the above plots. Paramater pairs that lead to the lowest cost (dark blue) in each terrain are highlighted in green circles.
@@ IV-D Model for Gait Parameter Generation
We pass the combined latent representation $\mathbf{Z}^{t}_{\text{comb}}$ through a multi-layer perceptron (MLP) regressor to predict suitable gait parameters ($\mathbf{g}$), specifically, the robot’s step height and hip splay. We train the regressor using the MSE loss, based on the training labels generated using the analysis of (5) on the dataset.
The regressor contains two fully connected hidden layers with ReLU activation, while the output layer does not use any activation function. ReLU introduces nonlinearity, enabling the network to model complex relationships between the combined representation and gait parameters.
The model learns continuous mappings from sensory inputs to gait parameters, allowing it to generalize to unseen terrains by interpolating or extrapolating from training patterns. This approach ensures dynamic gait adaptation based on real-time sensory data, providing a robust solution for navigating complex terrains. A dynamic window approach smoothens consecutive gait parameter predictions, ensuring stable and controlled gait transitions.
@@ V-A Implementation
We implemented CROSS-GAiT using PyTorch and trained all models on an NVIDIA A5000 GPU. For real-world experiments, we used the Ghost Vision 60 robot from Ghost Robotics, equipped with a front-facing wide-angle camera and an onboard Intel NUC 11 (Intel i7 CPU, NVIDIA RTX 2060 GPU). Our system runs in real-time at 60Hz. Path planning was implemented using [4]. We sample a 6-channel IMU and joint effort data, comprising 12 channels (one for each motor, measured in Nm), at 25 Hz. We combine these data streams to form the input time series data $\mathbf{X}^{t}_{\text{TS}}$, resulting in $n_{\text{ch}}=18$. $d_{\text{TS}}=100$, is chosen empirically to capture the frequency components of the robot’s cyclic gaits. $d_{m}=160$ based on [46].
@@
V-B Dataset Collection
Figure 5: Robot trajectories in unstructured terrains using CROSS-GAiT and comparison methods along with corresponding Joint Effort (left) and IMU Energy Density (right) plots are shown below each scenario. CROSS-GAiT reaches the goal faster with lower joint effort, while maintaining comparable IMU energy density to other methods.